3 years ago jane was one-half as old as she will be 2 years from now. what is janes age

Answer by richwmiller(17219)

j-3=1/2(j+2)
j-3=j/2+1
j/2=4
j=8
check
8-3=1/2(8+2)
5=1/2(10)
5=5
ok